OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and perpendicular to y =1/5x - 10. A:y = 5x + 10 B: y = 5x C:y = -5x - 10 D: y = -5x

OpenStudy (anonymous):

HELP!

OpenStudy (anonymous):

The line equation is y=kx+b, the lines are perpendicular, so k1*k2=-1, where k1=1/5, k2 is unknown. Solve for k2: k2=-5. So the line we are looking for is y=-5x+b. We know line is passing through the point (1,-5), which means -5=-5+b => b=0. The answer is D

OpenStudy (anonymous):

Thank you!

OpenStudy (mathstudent55):

From the given line, you can find its slope. Once you have the slope of the given line, you can find the slope of a perpendicular. Once you know the slope of the perpendicular, write y = mx + b, where m is the slope of the perpendicular. Plug in the given point (1, -5) and solve for b.